OpenStudy (anonymous):

ALGEBRA 1!!! MATH!! HELP!! I'M CLUELESS!

OpenStudy (lgbasallote):

syooor:DDD

OpenStudy (anonymous):

Simplify \[(2\sqrt{5} + 3\sqrt{7})^2\]

OpenStudy (anonymous):

The wait is killing me.

OpenStudy (anonymous):

use this ( a+b)^2 = a^2 +2ab + b^2

OpenStudy (anonymous):

Please explain steps and show your work! Thanks to all answers :D

OpenStudy (anonymous):

Use the foil rule. \[(2\sqrt{5}*3\sqrt{7})(2\sqrt{5}*3\sqrt{7})\]

OpenStudy (anonymous):

First, outter, inner, last.

OpenStudy (lgbasallote):

ugh...uhmmmm...i'm gonna use the a^2 + 2ab + b^2 thing \[(2\sqrt{5})^{2} + 2(2\sqrt{5})(3\sqrt{7}) + (3\sqrt{7})^{2}\]

OpenStudy (anonymous):

Whoops, instead of * replace with +

OpenStudy (anonymous):

I still don't understand!

OpenStudy (lgbasallote):

\[(2^{2})(\sqrt{5}^{2}) + (2)(2)(3)(\sqrt{(7)(5)} + (3^{2})(\sqrt{7}^{2})\] can you continue from here now? :D

OpenStudy (anonymous):